Harris Corp. manufactures a single product. Costs for the year 2001 for output levels of 1000 and 2000 units are as follows:

Units Produced 1000 2000
Direct labor $30,000 $30,000
Direct materials $20,000 $40,000
Overhead:
Variable portion $12,000 $24,000
Fixed portion $36,000 $36,000
Selling & Administrative costs:
Variable portion $5,000 $10,000
Fixed portion $22,000 $22,000

At each level of output, compute the following:
i. Total variable costs
ii. Total variable costs per unit
iii. Total costs that have to be recovered if the firm is to make a profit
1
Expert's answer
2012-06-21T11:08:06-0400
i. TVC1 = 20,000+12,000+5,000 = $37,000
& & TVC2 = 40,000+24,000+10,000 = $74,000
ii. TVC1 per unit = 67,000/1,000 = $37
& & TVC1 per unit = 104,000/2,000 = $37
iii. TC1 = $125,000
& & & TC2 = $162,000